From nch@roe.ac.uk Fri Aug 10 12:23:11 2012 Date: Fri, 10 Aug 2012 12:22:14 +0100 From: Nigel Hambly To: Eckhard Sutorius , Nicholas Cross , Rob Blake , Ross Collins , Mike Read , Mark Holliman , Clive Davenhall Cc: Lorenzo Rimoldini , Bob Mann , Jim Emerson , Keith Noddle , Norman Gray , Stephen Warren , Tom Shanks , Simon Dye , Stelios Voutsinas , Tom Kerr , Mike Irwin , Andy Lawrence at ROE Subject: WFAU VDFS Science Archives project team meeting minutes, 10/8/2012 Present: MAR, RSC, NJC, ETWS, MSH, KTN, NCH, RPB Apologies: JPE, AL, STV, RGM, NG DoNM: 10am Friday August 24th 2012 in the Vista Hut (subject to non-demolition) MAR in the chair. Actions discharged this time: ----------------------------- ACTION: NCH to nudge JRL and ask if VDFS-like bulk slurps of GES data can be set up. Discharged; JRL clearly very busy with GES pipeline processing issues given the flurry of emails, so NCH reluctant to keep badgering him... ACTION: NJC to communicate with Nottingham concerning incorporation of UDS in DR10. Discharged (by Omar!); looks like the timescales for production of the DR10 stacked mosaics and sorting out the final QC issues for the wide shallow surveys will converge, so a fabulous new UDS should appear as part of DR10 at first release. ACTION: NJC to contact VVV team to see if they're happy with world release of DR1. Discharged; clearly some rather worrying photometric issues identified (at the last minute) by Phil Lucas, which MAR and NJC will look into. ACTION: NJC to contact ESO and see if we could still send someone to the ESO workshop in October. Discharged; abstract submitted at the suggestion of the ESO LOC. Actions partly discharged but continuing: ----------------------------------------- ACTION: RSC to scope out the areas/situations where we'd like to test MonetDB i.e. where do the current bottlenecks lie? Continues; ACTION: ETWS (with help from NJC as necessary) to add FeII 1.644 micron filter details into the WSA. Continues; ETWS had made some code changes but further set-up required. Actions carried forward from 27/07/12 meeting --------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. Specific points and new actions: -------------------------------- Project management: Nothing to report this time. WFCAM, VISTA and VST updates: Nothing to report this time. Comments and issues arising from CASU minutes: The team noted the minutes of the meeting of 25 July; there were no comments, other than there was a disappointing lack of mots franglais cette fois. Networking: Network transfers continue to work well. ETWS noted: WSA/VSA/OSA/GES Operations: ETWS noted: -- Catalogue ingests have been done up till May 2012 for the WSA, and up till March 2012 for the VSA. The VVV catalogue ingests are just running for September 2011. -- OSA metadata has been ingested completely up till June 2012. -- Ran the ESO Release code for VVVDR1 and VMC. -- Finalised the OSA development branch and merged it back into trunk. -- Updated the WSA monitor pages for the UHS survey. Hardware and Systems: All quiet on the hardware front. Survey Data Release: UKIDSS DR10 awaits the final QC results for pawsky deprecations in the wide shallow surveys ACTION: NCH to check the GCS for pawsky-related issues, and liaise with SJW over the LAS results and finalisation of the phase-1 QC deprecations before generation of the merged source tables. Looks like the new UDS stacked mosaics can/will be delivered on the same timescale. Vista catalogue phase-3 deliverables are pretty much done now, and simply need to be uploaded, except for VIKING which requires list-driven photometry (but last year's image phase-2 deliverables have been done - no feedback from ESO so far). NJC noted that VMC requirements have shifted a little in that some additional ancillary data ("PSF tables") are to be included. New survey team consortium releases, including data up to March this year, are now being generated in the order VMC, VHS, VIKING. Software: RSC noted: "GES archive's atomic line database schema has been revised allowing the complete set of version 2 data to be ingested. A plan has been proposed for modifying the detection quality flagging algorithm of saturated sources in VISTA tiles. Using the average saturation level value of the detector frame, as we do for the other frame types, was unsuitable for tiles due to the large degree of spatial variation resulting in too many saturated sources not being flagged and thus leading to spurious variable sources, which the VVV in particular are concerned about. We will now flag saturated sources from VISTA tiles based on the minimum average saturation level of the component frames of the tile, which should avoid the spurious variables problem. We also realised that the detection quality flagging process will have to be updated for the VVV to support the new monthly-split detection table schema, whereby each individual table will need to be updated since the unifying view is read-only." NJC noted that (hopefully) final mods have been made to the ESO deliverables software, with many small fixes required to accomodate ESO's and the PI's requirements. MAR noted that he has made some UI mods to enable easier UHS QC; finally ETWS noted that the VDFS "OSA branch" software has been merged back into the SVN trunk, such that we now have a software suite that works for all of VSA, WSA and OSA for Vista, WFCAM and VST. All expressed surprise and delight that this hasn't totally knackered everything. Non-survey Data Release: A new release of w/serv/6 has been made for the PI. Astrogrid deployment & Data Analysis services: MSH reported continuing work on the Astrogrid CVS and Wiki services being moved to ROE; important core services are safely migrated and running at WFAU now. KTN noted that STV is working away tenaciously on the new web query interface, with integration into SAMP and OGSA-DAI federated dataset querying. The plan is to deploy on the first release of VST-ATLAS in the OSA. Miscellaneous: This year's ROE Workshop has been announced: "Synergies between large-area infrared surveys, VLT-MOONS and Euclid" http://www.roe.ac.uk/roe/workshop/2012/ With loads of overlap with the WFAU projects a good presence would be useful. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. ============================================================= -- Scanned by iCritical.